# Sparse juntas on the biased hypercube

### v4

@@ 2 Preliminaries
Our basic notion of closeness is $L_{2}$. Two functions $f,g$ are $\epsilon$-close with respect to $\mu_{p}$ if $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]\leq\epsilon$. If $\mu_{p}$ is clear from context then we omit its mention.
@@
A function $f\colon\{0,1\}^{n}\to\mathbb{R}$ has degree $d$ if $\deg f\leq d$. A function is $\epsilon$-close to degree $d$ (with respect to $\mu_{p}$) if $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]\leq\epsilon$ for some degree $d$ function $g$. Equivalently, $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f^{>d})^{2}]\leq\epsilon$, where $f^{\leq d},f^{>d}$ are obtained from the $p$-biased Fourier expansion of $f$ (described in Section 4.1) by retaining all levels up to $d$ and all levels beyond $d$, respectively.
@@
A function is Boolean if it is $\{0,1\}$-valued. If $f,g$ are Boolean then they are $\epsilon$-close iff $\Pr[f\neq g]\leq\epsilon$. In other words, for Boolean functions the $L_{2}$ and $L_{0}$ notions of distance coincide.
If $f\colon\{0,1\}^{n}\to\mathbb{R}$ then $\operatorname{round}(f,\{0,1\})$ is the Boolean function obtained by rounding each $f(y)$ to the nearest value among $\{0,1\}$. The original function $f$ is $\epsilon$-close to Boolean if
$\operatorname*{\mathbb{E}}_{y}[\operatorname{dist}(f(y),\{0,1\})^{2}]\leq\epsilon,$
where $\operatorname{dist}(f(y),\{0,1\})=\min(|f(y)-0|,|f(y)-1|)$. Equivalently, $\operatorname*{\mathbb{E}}[(f-\operatorname{round}(f,\{0,1\}))^{2}]\leq\epsilon$.
@@ 2.1 Unbiased structure theorems
The claim follows since $2^{|S|-1}$ of the summands are $+f(z)$ and $2^{|S|-1}$ are $-f(z)$. ∎
@@
Nisan and Szegedy [NS94] proved the following theorem.
@@
The number of coordinates in the junta was improved to $O(2^{d})$ in [CHS20, Wel22].
@@
In unpublished work, Kindler and Safra [KS02, Kin03] proved the following result, which we state in its formulation due to Keller and Klein [KK20].
@@
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\epsilon$-close to degree $d$, where $\epsilon<c^{d}$, then $f$ is $2\epsilon$-close to a Boolean degree $d$ function.
This formulation follows by combining Theorem 1.4 and Proposition 5.6 in [KK20]. (Keller and Klein also prove a stronger version in which the closeness is improved to the optimal expression, $\epsilon+\tilde{O}(\epsilon^{2})$.)
For our purposes, we need a version of the Kindler–Safra theorem which holds for all $\epsilon$.
@@
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\epsilon$-close to degree $d$ then $f$ is $C^{d}\epsilon$-close to a Boolean degree $d$ function.
Proof.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\epsilon$-close to degree $d$. If $\epsilon<c^{d}$ then the result follows directly from Theorem 2.3. Otherwise, $C^{d}\epsilon\geq 1$, and so $f$ is $C^{d}\epsilon$-close to the constant zero function. ∎
@@
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$ then $f$ is $K_{d}\epsilon$-close to a monotone Boolean degree $d$ function.
Proof.
Suppose that $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is monotone and $\epsilon$-close to degree $d$. Applying Theorem 2.4, $f$ is $C^{d}\epsilon$-close to a Boolean degree $d$ function $g$.
@@
On the other hand, this probability is at most $C^{d}\epsilon$, and so $\epsilon\geq C^{-d}/2^{2d-1}$. It follows that $f$ is $2^{2d-1}C^{d}\epsilon$-close to the zero function. ∎
@@ 2.2 Forbidden configurations
Our work involves three different types of forbidden configurations. These are configurations of coefficients which cannot occur in Boolean functions, but do arise in our context — but only sparingly. For example, the biased FKN theorem [Fil16] involves functions of the form $f=\sum_{i}c_{i}y_{i}$. If $f$ is Boolean then we cannot have $c_{i}=c_{j}=1$, but if $f$ is $\epsilon$-close to Boolean then there could be up to $O(\epsilon/p^{2})$ copies of this configuration.
@@
Proof.
Let $g=f(f-1)$, and observe that $x$ is a minimal non-Boolean input of $f$ iff it is a minimal non-zero input of $g$, that is, $g(x)\neq 0$ but $g(y)=0$ for all $y<x$. If $\mathbbm{1}_{S}$ is a minimal non-zero input of $g$ then $\tilde{g}(S)\neq 0$, and so $|S|\leq\deg g\leq 2d$. Moreover, $S$ can be written as the union of at most two sets in $\operatorname{supp}(f)$. ∎
@@
Proof.
By minimality, if $\mathbbm{1}_{S}$ is a minimal non-monotone input of $f$ then $f|_{S}$ depends on all inputs, and so $|S|\leq M_{d}$ by Theorem 2.2. Moreover, $S$ is the union of all sets in $\operatorname{supp}(f|_{S})$, and so it is the union of at most $|S|$ sets in $\operatorname{supp}(f)$. ∎
@@
If $|S|\leq M_{d}$ then we are done. Otherwise, let $m$ be the minimal index such that $|B_{1}\cup\cdots\cup B_{m}|>M_{d}$. By minimality, $\deg(f|_{B_{1}\cup\cdots\cup B_{m-1}})\leq d$, and so $|B_{1}\cup\cdots\cup B_{m-1}|\leq M_{d}$ by Theorem 2.2. Since $f|_{B_{1}\cup\cdots\cup B_{m}}$ depends on all inputs and $|B_{1}\cup\cdots\cup B_{m}|>M_{d}$, Theorem 2.2 shows that it cannot have degree $d$, and so $B_{1}\cup\cdots\cup B_{m}=S$. Therefore $|S|\leq M_{d}+d$. ∎
@@ 3 Sparsity
A set system $\mathcal{F}$ is $(d,C,\epsilon)$-sparse, where $d\geq 0$ is an integer, $\epsilon\geq 0$, and $C\geq 1$, if the following three properties hold:
@@
    For all integer $e\geq 0$, the set system $\mathcal{F}$ contains at most $C^{e}\epsilon$ sets $U$ of size $e$.
@@
We describe some elementary properties of sparse set systems in Section 3.1. Sparse set systems satisfy a crucial property we call the reverse union bound, which we explain in Section 3.2. Finally, we discuss quantized sparse juntas, which are sparse juntas in which the monomial coefficients belong to a fixed finite set. We show in Section 3.3 that if a quantized sparse junta $f$ is $L_{0}$-close to Boolean (that is, $\Pr[f\notin\{0,1\}]\leq\epsilon$) then it is also $L_{2}$-close to Boolean (that is, $\operatorname*{\mathbb{E}}[\operatorname{dist}(f,\{0,1\})^{2}]=O(\epsilon)$).
@@ 3.1 Elementary properties
Let $\mathcal{F},\mathcal{G}$ be $(d,C,\epsilon)$-sparse.
@@
    $\mathcal{F}\cup\mathcal{G}$ is $(d,2C,\epsilon)$-sparse.
@@
    $\mathcal{F}\sqcup\mathcal{G}$ is $(2d,O_{d}(C),\epsilon)$-sparse. For this it suffices that $\mathcal{G}$ be $(d,C)$-sparse.
@@
    $\mathcal{F}^{\sqcup k}$ are $(kd,O_{k,d}(C),\epsilon)$-sparse.
Proof.
@@
Let $\mathcal{F},\mathcal{G}$ be $(d,C,\epsilon)$-sparse set systems. Lemma 3.5 shows that $\mathcal{F}\cup\mathcal{G}$ is $(d,2C)$-sparse. Given $e\geq 1$, there are at most $C^{e}\epsilon$ sets of size $e$ in each of $\mathcal{F},\mathcal{G}$. Therefore there are at most $2C^{e}\epsilon\leq(2C)^{e}\epsilon$ sets of size $e$ in $\mathcal{F}\cup\mathcal{G}$.
The case $e=0$ requires special treatment. If $\epsilon<1$ then $\mathcal{F},\mathcal{G}$ do not contain $\emptyset$, and so $\mathcal{F}\cup\mathcal{G}$ also doesn’t contain $\emptyset$. If $\epsilon\geq 1$ then clearly $\mathcal{F}\cup\mathcal{G}$ contains at most $\epsilon$ sets of size 00.
@@
Let $\mathcal{F}$ be a $(d,C,\epsilon)$-sparse set system and let $\mathcal{G}$ be a $(d,C)$-sparse set system. Lemma 3.5 shows that $\mathcal{F}\sqcup\mathcal{G}$ is $(d,O_{d}(C))$-sparse.
Given $e$, we need to bound the number of sets $S=A\cup B$ of size $e$, where $A\in\mathcal{F}$ and $B\in\mathcal{G}$. For this, we show that for each $a,a_{b}\leq d$, there are at most $2^{d}C^{e}\epsilon$ sets $A\cup B$ of size $e$ such that $A\in\mathcal{F}$, $B\in\mathcal{G}$, $|A|=a$, and $|A\cap B|=a_{b}$.
Indeed, there are at most $C^{a}\epsilon$ sets $A\in\mathcal{F}$ of size $a$. Given $A$, there are at most $\binom{|A|}{a_{b}}\leq 2^{d}$ choices for $A\cap B$. Since $|A\cup B|=e$, we have $|B|=|A\cup B|+|A\cap B|-|A|=e+a_{b}-a$. Therefore given $A\cap B$, there are at most $C^{e-a}$ sets $B\in\mathcal{G}$ including $A\cap B$ such that $|A\cup B|=e$. Altogether, given $a,a_{b}$, the number of pairs $A,B$ satisfying the properties listed above is bounded by
$\underbrace{C^{a}\epsilon}_{A}\cdot\underbrace{2^{d}}_{A\cap B\text{ given }A}\cdot\underbrace{C^{e-a}}_{B\text{ given }A\cap B}=2^{d}C^{e}\epsilon,$
@@
Since there are at most $(d+1)^{2}$ many options for $a,a_{b}$, this shows that the number of sets $S\in\mathcal{F}\sqcup\mathcal{G}$ of size $e\geq 1$ is at most $(d+1)^{2}2^{d}C^{e}\epsilon\leq((d+1)^{2}2^{d}C)^{e}\epsilon$, completing the proof except for the case $e=0$.
In order to handle the case $e=0$, we again consider whether $\epsilon<1$ or not. If $\epsilon<1$ then $\emptyset\notin\mathcal{F}$ and so $\emptyset\notin\mathcal{F}\sqcup\mathcal{G}$. If $\epsilon\geq 1$ then $\mathcal{F}\sqcup\mathcal{G}$ trivially contains at most $\epsilon$ sets of size 00.
@@
This item follows from Item b by induction on $k$. ∎
@@ 3.2 Reverse union bound
In the other direction, observe first that the events $\mathcal{E}_{S}$ are disjoint. Indeed, suppose that $S,T\in\mathcal{F}$ are distinct, and assume without loss of generality that $S\not\subseteq T$. If $\mathcal{E}_{S}$ holds then $y_{S}=1$ and so $\mathcal{E}_{T}$ cannot hold. This implies that
@@
where (*)(*)( * ) follows from the Harris–FKG inequality, since the events $y_{T}=0$ are anti-monotone. Lemma 3.5 shows that $\mathcal{G}|_{S\leftarrow 1}$ is $(d,K/p)$-sparse, where $K$ depends only on $d,C$. Therefore
@@
$\Pr[\mathcal{E}]=\sum_{S\in\mathcal{F}}\Pr[\mathcal{E}_{S}]=\Omega_{d,C}\left(\sum_{S\in\mathcal{F}}p^{|S|}\right).\qed$

### v6

@@ 2 Preliminaries
Our basic notion of closeness is $L_{2}$. Two functions $f,g$ are $\varepsilon$-close with respect to $\mu_{p}$ if $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]\leq\varepsilon$. If $\mu_{p}$ is clear from context then we omit its mention.
@@
A function $f\colon\{0,1\}^{n}\to\mathbb{R}$ has degree $d$ if $\deg f\leq d$. A function is $\varepsilon$-close to degree $d$ (with respect to $\mu_{p}$) if $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-g)^{2}]\leq\varepsilon$ for some degree $d$ function $g$.
We can also define these concepts via the $p$-biased Fourier expansion. Let $f=\sum_{S}\hat{f}(S)\omega_{S}$, where $\{\omega_{S}\}$ is the $p$-biased Fourier basis, described in Section 4.1. Let $f^{\leq d}=\sum_{|S|\leq d}\hat{f}(S)\omega_{S}$ and $f^{>d}=\sum_{|S|>d}\hat{f}(S)\omega_{S}$. The function $f$ has degree $d$ if $f^{>d}=0$. The degree $d$ function closest to $f$ (with respect to $\mu_{p}$) is $f^{\leq d}$, and so $f$ is $\varepsilon$-close to degree $d$ if $\operatorname*{\mathbb{E}}_{\mu_{p}}[(f^{>d})^{2}]=\operatorname*{\mathbb{E}}_{\mu_{p}}[(f-f^{\leq d})^{2}]\leq\varepsilon$.
@@
A function is Boolean if it is $\{0,1\}$-valued. If $f,g$ are Boolean then they are $\varepsilon$-close if and only if $\Pr[f\neq g]\leq\varepsilon$. In other words, for Boolean functions the $L_{2}$ and $L_{0}$ notions of distance coincide.
If $f\colon\{0,1\}^{n}\to\mathbb{R}$ then $\operatorname{round}(f,\{0,1\})$ is the Boolean function obtained by rounding each $f(y)$ to the nearest value among $\{0,1\}$. The original function $f$ is $\varepsilon$-close to Boolean if
$\operatorname*{\mathbb{E}}_{y}[\operatorname{dist}(f(y),\{0,1\})^{2}]\leq\varepsilon,$
where $\operatorname{dist}(f(y),\{0,1\})=\min(|f(y)-0|,|f(y)-1|)$.
Equivalently, $\operatorname*{\mathbb{E}}[(f-\operatorname{round}(f,\{0,1\}))^{2}]\leq\varepsilon$.
@@ 2.1 Unbiased structure theorems
The claim follows since $2^{|S|-1}$ of the summands are $+f(z)$ and $2^{|S|-1}$ are $-f(z)$.
@@
Nisan and Szegedy [NisanS1994] proved the following theorem.
@@
The number of coordinates in the junta was improved to $O(2^{d})$ in [ChiarelliHS2020, Wellens2022].
@@
In unpublished work, Kindler and Safra [KindlerS2002, Kindler2003] proved the following result, which we state in its formulation due to Keller and Klein [KellerK2020].
@@
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\varepsilon$-close to degree $d$, where $\varepsilon<c^{d}$, then $f$ is $2\varepsilon$-close to a Boolean degree $d$ function.
This formulation follows by combining Theorem 1.4 and Proposition 5.6 in [KellerK2020]. (Keller and Klein also prove a stronger version in which the closeness is improved to the optimal expression, $\varepsilon+\tilde{O}(\varepsilon^{2})$.)
For our purposes, we need a version of the Kindler–Safra theorem which holds for all $\varepsilon$.
@@
There exists a universal constant $C>0$ such that the following holds for all $d\in\mathbb{N}$. If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\varepsilon$-close to degree $d$ then $f$ is $C^{d}\varepsilon$-close to a Boolean degree $d$ function.
Proof 2.6.
@@
Suppose that $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is $\varepsilon$-close to degree $d$. If $\varepsilon<c^{d}$ then the result follows directly from Theorem 2.4. Otherwise, $C^{d}\varepsilon\geq 1$, and so $f$ is $C^{d}\varepsilon$-close to the constant zero function.
@@
If $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$ then $f$ is $K_{d}\varepsilon$-close to a monotone Boolean degree $d$ function.
Proof 2.10.
Suppose that $f\colon(\{0,1\}^{n},\mu_{1/2})\to\{0,1\}$ is monotone and $\varepsilon$-close to degree $d$. Applying Theorem 2.5, $f$ is $C^{d}\varepsilon$-close to a Boolean degree $d$ function $g$.
@@
On the other hand, this probability is at most $C^{d}\varepsilon$, and so $\varepsilon\geq C^{-d}/2^{2d-1}$. It follows that $f$ is $2^{2d-1}C^{d}\varepsilon$-close to the zero function.
@@ 2.2 Forbidden configurations
Our work involves three different types of forbidden configurations. These are configurations of coefficients which cannot occur in Boolean functions, but do arise in our context — but only sparingly. For example, the biased FKN theorem [Filmus2016-fkn] involves functions of the form $f=\sum_{i}c_{i}y_{i}$. If $f$ is Boolean then we cannot have $c_{i}=c_{j}=1$, but if $f$ is $\varepsilon$-close to Boolean then there could be up to $O(\varepsilon/p^{2})$ copies of this configuration.
@@
Proof 2.12.
Let $g=f(f-1)$, and observe that $x$ is a minimal non-Boolean input of $f$ if and only if it is a minimal non-zero input of $g$, that is, $g(x)\neq 0$ but $g(y)=0$ for all $y<x$. If $\mathbbm{1}_{S}$ is a minimal non-zero input of $g$ then $\tilde{g}(S)\neq 0$, and so $|S|\leq\deg g\leq 2d$. Moreover, $S$ can be written as the union of at most two sets in $\operatorname{supp}(f)$.
@@
Proof 2.14.
By minimality, if $\mathbbm{1}_{S}$ is a minimal non-monotone input of $f$ then $f|_{S}$ depends on all inputs, and so $|S|\leq M_{d}$ by Theorem 2.3. Moreover, $S$ is the union of all sets in $\operatorname{supp}(f|_{S})$, and so it is the union of at most $|S|$ sets in $\operatorname{supp}(f)$.
@@
If $|S|\leq M_{d}$ then we are done. Otherwise, let $m$ be the minimal index such that $|B_{1}\cup\cdots\cup B_{m}|>M_{d}$. By minimality, $\deg(f|_{B_{1}\cup\cdots\cup B_{m-1}})\leq d$, and so $|B_{1}\cup\cdots\cup B_{m-1}|\leq M_{d}$ by Theorem 2.3. Since $f|_{B_{1}\cup\cdots\cup B_{m}}$ depends on all inputs and $|B_{1}\cup\cdots\cup B_{m}|>M_{d}$, Theorem 2.3 shows that it cannot have degree $d$, and so $B_{1}\cup\cdots\cup B_{m}=S$. Therefore $|S|\leq M_{d}+d$.
@@ 3 Sparsity
A set system $\mathcal{F}$ is $(d,C,\varepsilon)$-sparse, where $d\geq 0$ is an integer, $\varepsilon\geq 0$, and $C\geq 1$, if the following three properties hold:
@@
    For all integer $e\geq 0$, the set system $\mathcal{F}$ contains at most $C^{e}\varepsilon$ sets $U$ of size $e$.
@@
We describe some elementary properties of sparse set systems in Section 3.1. Sparse set systems satisfy a crucial property we call the reverse union bound, which we explain in Section 3.2. Finally, we discuss quantized sparse juntas, which are sparse juntas in which the monomial coefficients belong to a fixed finite set. We show in Section 3.3 that if a quantized sparse junta $f$ is $L_{0}$-close to Boolean (that is, $\Pr[f\notin\{0,1\}]\leq\varepsilon$) then it is also $L_{2}$-close to Boolean (that is, $\operatorname*{\mathbb{E}}[\operatorname{dist}(f,\{0,1\})^{2}]=O(\varepsilon)$).
@@ 3.1 Elementary properties
Let $\mathcal{F},\mathcal{G}$ be $(d,C,\varepsilon)$-sparse.
@@
    $\mathcal{F}\cup\mathcal{G}$ is $(d,2C,\varepsilon)$-sparse.
@@
    $\mathcal{F}\sqcup\mathcal{G}$ is $(2d,O_{d}(C),\varepsilon)$-sparse. For this it suffices that $\mathcal{G}$ be $(d,C)$-sparse.
@@
    $\mathcal{F}^{\sqcup k}$ are $(kd,O_{k,d}(C),\varepsilon)$-sparse.
Proof 3.8.
@@
Let $\mathcal{F},\mathcal{G}$ be $(d,C,\varepsilon)$-sparse set systems. Lemma 3.5 shows that $\mathcal{F}\cup\mathcal{G}$ is $(d,2C)$-sparse. Given $e\geq 1$, there are at most $C^{e}\varepsilon$ sets of size $e$ in each of $\mathcal{F},\mathcal{G}$. Therefore there are at most $2C^{e}\varepsilon\leq(2C)^{e}\varepsilon$ sets of size $e$ in $\mathcal{F}\cup\mathcal{G}$.
The case $e=0$ requires special treatment. If $\varepsilon<1$ then $\mathcal{F},\mathcal{G}$ do not contain $\emptyset$, and so $\mathcal{F}\cup\mathcal{G}$ also doesn’t contain $\emptyset$. If $\varepsilon\geq 1$ then clearly $\mathcal{F}\cup\mathcal{G}$ contains at most $\varepsilon$ sets of size 00.
@@
Let $\mathcal{F}$ be a $(d,C,\varepsilon)$-sparse set system and let $\mathcal{G}$ be a $(d,C)$-sparse set system. Lemma 3.5 shows that $\mathcal{F}\sqcup\mathcal{G}$ is $(d,O_{d}(C))$-sparse.
Given $e$, we need to bound the number of sets $S=A\cup B$ of size $e$, where $A\in\mathcal{F}$ and $B\in\mathcal{G}$. For this, we show that for each $a,a_{b}\leq d$, there are at most $2^{d}C^{e}\varepsilon$ sets $A\cup B$ of size $e$ such that $A\in\mathcal{F}$, $B\in\mathcal{G}$, $|A|=a$, and $|A\cap B|=a_{b}$.
Indeed, there are at most $C^{a}\varepsilon$ sets $A\in\mathcal{F}$ of size $a$. Given $A$, there are at most $\binom{|A|}{a_{b}}\leq 2^{d}$ choices for $A\cap B$. Since $|A\cup B|=e$, we have $|B|=|A\cup B|+|A\cap B|-|A|=e+a_{b}-a$. Therefore given $A\cap B$, there are at most $C^{e-a}$ sets $B\in\mathcal{G}$ including $A\cap B$ such that $|A\cup B|=e$. Altogether, given $a,a_{b}$, the number of pairs $A,B$ satisfying the properties listed above is bounded by
$\underbrace{C^{a}\varepsilon}_{A}\cdot\underbrace{2^{d}}_{A\cap B\text{ given }A}\cdot\underbrace{C^{e-a}}_{B\text{ given }A\cap B}=2^{d}C^{e}\varepsilon,$
@@
Since there are at most $(d+1)^{2}$ many options for $a,a_{b}$, this shows that the number of sets $S\in\mathcal{F}\sqcup\mathcal{G}$ of size $e\geq 1$ is at most $(d+1)^{2}2^{d}C^{e}\varepsilon\leq((d+1)^{2}2^{d}C)^{e}\varepsilon$, completing the proof except for the case $e=0$.
In order to handle the case $e=0$, we again consider whether $\varepsilon<1$ or not. If $\varepsilon<1$ then $\emptyset\notin\mathcal{F}$ and so $\emptyset\notin\mathcal{F}\sqcup\mathcal{G}$. If $\varepsilon\geq 1$ then $\mathcal{F}\sqcup\mathcal{G}$ trivially contains at most $\varepsilon$ sets of size 00.
@@
This item follows from Item b by induction on $k$.
@@ 3.2 Reverse union bound
In the other direction, observe first that the events $\mathcal{E}_{S}$ are disjoint. Indeed, suppose that $S,T\in\mathcal{F}$ are distinct. If $S\subseteq T$ and $T\subseteq S$ then $S=T$, so assume without loss of generality that $S\not\subseteq T$. If $\mathcal{E}_{S}$ holds then $y_{S}=1$ and so $\mathcal{E}_{T}$ cannot hold. This implies that
@@
where ()(*)( ∗ ) follows from the Harris–FKG inequality, since the events $y_{T}=0$ are anti-monotone. Lemma 3.5 shows that $\mathcal{G}|_{S\leftarrow 1}$ is $(d,K/p)$-sparse, where $K$ depends only on $d,C$. Therefore
@@
$\Pr[\mathcal{E}]=\sum_{S\in\mathcal{F}}\Pr[\mathcal{E}_{S}]=\Omega_{d,C}\left(\sum_{S\in\mathcal{F}}p^{|S|}\right).$